Prove that Cot 2A +tan A= cosec 2A.. Can you please answer this question immediately.. It`s urgent

Grade:11

1 Answers

Arun
6 years ago
cot 2A + tan A 

= cos 2A/sin 2A + sin A/cos A 

= (cos 2A cos A + sin 2A sin A) / sin 2A cos A 

Using the compound angle formulae, 

we know that cos (2A - A ) = cos 2A cos A + sin 2A sin A... 

therefore, 

= cos (2A - A) / sin 2A cos A 
= cos A / sin 2A cos A 
= 1/ sin 2A 
= cosec 2A